OpenStudy (anonymous):

PLEASE HELP! please I need the work shown. Algebra 1 Word Problem "You have a part-time job. You work for 3 hrs on Friday and 6 hours on Saturday. You also receive allowance of $20 per week. you earn $92 per week. How much do you earn per hour at your part-time job" ?

OpenStudy (anonymous):

92 - 20 = 72 72/9 = 8

OpenStudy (anonymous):

92-20=72 72/9= 8 Answer is 8

OpenStudy (anonymous):

THANK YOU BOTH SO SO MUCH!!!